OpenStudy (anonymous):

Urn A has 16 white and 10 red balls. Urn B has 2 white and 9 red balls. We flip a fair coin. If the outcome is heads, then a ball from urn A is selected, whereas if the outcome is tails, then a ball from urn B is selected. Suppose that a red ball is selected. What is the probability that the coin landed heads?

OpenStudy (anonymous):

@is3535

OpenStudy (is3535):

9 didve by 2=

OpenStudy (anonymous):

how'd you get 9/2? @is3535

OpenStudy (is3535):

Urn B has 2 white and 9 red balls

OpenStudy (anonymous):

9/2 is 4.5 @is3535

OpenStudy (is3535):

that your answer

OpenStudy (anonymous):

4.5?

OpenStudy (anonymous):

4.5? its asking the probability? @is3535

OpenStudy (anonymous):

@Leong @YoloShroom

OpenStudy (is3535):

oh im sry it will be 9

OpenStudy (is3535):

whic is biger 9 or 2

OpenStudy (anonymous):

9

OpenStudy (is3535):

yea

OpenStudy (is3535):

9 is your answer

OpenStudy (anonymous):

how? @is3535

OpenStudy (is3535):

Urn B has 2 white and 9 red balls. We flip a fair coin. If the outcome is heads, then a ball from urn A is selected, whereas if the outcome is tails, then a ball from urn B is selected. Suppose that a red ball is selected. What is the probability that the coin landed head,

OpenStudy (anonymous):

are we doing some ratio??? or like....

OpenStudy (is3535):



OpenStudy (anonymous):

its just saying, Suppose that a red ball is selected. What is the probability that the coin landed heads? @Leong

OpenStudy (anonymous):

4.5 :v you take 9 divide for 2, like you got less ball mean less chance.

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Shalante

OpenStudy (pulsified333):

Probability can not be more than 1

OpenStudy (anonymous):

okay :v then if not one we should do some ratio here: 10 over 16, means that 0.625 chance that the A get to choose :v

OpenStudy (anonymous):

i don't think thats right @Leong

OpenStudy (anonymous):

@BAdhi

OpenStudy (badhi):

This involves conditional probability and total probability theorem @vzforever are you familiar with them?

OpenStudy (anonymous):

yes @BAdhi

OpenStudy (badhi):

can you show me the given information written as the notation used in probability such as p(A) , p(B) etc

OpenStudy (anonymous):

i don't understand @BAdhi

OpenStudy (badhi):

can you write down the given information in the mathematical notation? for example probability of choosing A => P(A)

OpenStudy (anonymous):

Pr(A)= 1/2 Pr(B)= 1/2 @BAdhi

OpenStudy (anonymous):

what next? @BAdhi

OpenStudy (badhi):

what about the probability of getting a red ball given that we have choosen A? can you show it in the notation?

OpenStudy (anonymous):

Pr(AR)=10/26

OpenStudy (badhi):

umm, by P(AR) you mean \(P(A \cap R)\) ?

OpenStudy (anonymous):

no its 10/42 @BAdhi

OpenStudy (badhi):

10/42 is the probability of choosing a red ball out of all the balls.. but you see, since its already given that A has been choosen the amount of balls are reduced to 26. And i still need the answer to the previous question @vzforever

OpenStudy (anonymous):

whats ur previous question? @BAdhi

OpenStudy (badhi):

what do you mean by P(AR).. from what ive learned no such notation is used in prbability either it has to be \(P(A\cap R) , P(A\cup R), P(A|R)\) ??

OpenStudy (anonymous):

yes ur right

OpenStudy (anonymous):

what u said before was correct

OpenStudy (badhi):

ok.. what about the notation.. arent you going to answer my question?

OpenStudy (anonymous):

i said yes its (P(A \cap R)\)

OpenStudy (badhi):

\(P(A\cap R)\) means the probility of choosing a red ball and choosing A in here event -> choosing A and choosing red ball has still not occured but what i asked was, probability of choosing a Red ball given that A is choosen. in here A has already being chosen

OpenStudy (badhi):

so the correct notation is \(P(R | A)\)

OpenStudy (anonymous):

ok

OpenStudy (badhi):

similarly you can find P(B), P(R|B), P(W|B), P(W|A) so since the event -> turning head and the event -> choosing A are both same, what they are asking is, P(A|R)

OpenStudy (anonymous):

ok so is it (1/2)/(10/42)

OpenStudy (anonymous):

which is 5/42

OpenStudy (badhi):

so what you are saying is, \(P(A|R) = P(R)\times P(A)\) ??

OpenStudy (anonymous):

can you explain to me? i need to get this done.

OpenStudy (anonymous):

@kropot72

OpenStudy (badhi):

Its normal to get confused with the difference between \(P(A\cap R)\) and \(P(R|A)\) so I recommend you to read and look more into the explanation ive given in the previous post The definition of the conditional probability is, \(P(R|A) = \frac{P(R\cap A)}{P(A)}\) since P(R|A) and P(A) are known, you can find \(P(R\cap A)\) since what they are asking is P(A|R) , \(P(A|R) =\frac{P(A\cap R)}{P(R)} \) can \(P(A\cap R)\) is obtained in the previous step and hope you know how to find P(R)

OpenStudy (anonymous):

ok so what do i do to solve this problem

OpenStudy (anonymous):

we haven't even gotten there yet

OpenStudy (anonymous):

@TQKMB

OpenStudy (anonymous):

@BAdhi can we do the problem?

OpenStudy (badhi):

thats what im trying to do the whole time.. i dont wanna give you the answer straight forward, since itll be pretty useless. as a guidance let me state this, you have to find the values of the following variables from the given information straight forward P(A), P(R|A), P(R) with use of conditional probability definition (which ive given above) you have to find \(P(A\cap R)\) after that find P(A|R) with use of the same conditional probability definition (which is also given previously) what they are asking is the value of P(A|R)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can u explain with numbers

OpenStudy (badhi):

it is really important to know the difference between P(A|R) and \(P(A \cap R) \) since , as you can see even in this problem they are applied so close together in problems Are you clear with the difference between them?

OpenStudy (badhi):

first give me what youve tried pls after that ill give the answer

OpenStudy (anonymous):

110/299 and 299/462 are the answers I've gotten and neither are right. obviously i need help cause idk what I'm doing. and I've made a tree

OpenStudy (anonymous):

@BAdhi

OpenStudy (anonymous):

@dan815